OpenStudy (anonymous):

Please help: 6x^3+7x^2-6x-7 can anyone help me solve for x?

OpenStudy (anonymous):

6x^3+7x^2-6x-7 = (6x^3-6x)+ (7x^2 -7) just regrouping = 6x( x^2 -1) +7( x^2 -1) factoring 6x ...and 7 = ( x^2 -1)(6x+7) factoring ( x^2 -1) now ( x^2 -1) =0 or 6x+7=0 x^2 = 1 or 6x= -7 x = +1 or x = -1 or x=-7/6